OpenStudy (anonymous):

Use the Change of Base Formula to evaluate log(8)77

OpenStudy (anonymous):

These are my choices... 3.467 1.886 4.344 2.089

jimthompson5910 (jim_thompson5910):

Similar Problem: \[\Large \log_{5}(32) = \frac{\log(32)}{\log(5)}\] \[\Large \log_{5}(32) = \frac{1.505149978}{0.6989700043}\] \[\Large \log_{5}(32) = 2.153382790\]

OpenStudy (anonymous):

so it would be 2.0889 ?

jimthompson5910 (jim_thompson5910):

good, \[\Large \log_{8}(77) = \frac{\log(77)}{\log(8)}\] \[\Large \log_{8}(77) = \frac{1.88649072517249}{0.90308998699194}\] \[\Large \log_{8}(77) = 2.0889288468983\]

OpenStudy (anonymous):

Remember the formula\[\log_a(M)=\frac{\log_b(M)}{\log_b(a)}\]So\[\log_8(77)=\frac{\log77}{\log8}=\frac{\ln77}{\ln8}\]

OpenStudy (anonymous):

(: thanks

OpenStudy (anonymous):

For a simpler way to solve, consider that \[8^2=64<77<8^3=512 \implies 2<\log_877<3\]which leads you directly to only one of your possible answers.